A new robot could be acquired for 60,000, and be kept for up to 4 years. Its annual expenses would be $6,000, $9,000, $14,000, and $20.000 in years 1 to 4 in the same order. The robot's market value would be $45,000, $33,750, $25,500 and $19,500 at the end of years 1 to 4 in the same order. Determine the economic life of the robot, assuming 10% cost of capital.
Answer #1

Year 1:
=-60000-6000/1.1+45000/1.1=-24545.4545454546

Year 2:
=-60000-6000/1.1-9000/1.1^2+33750/1.1^2=-45000

Year 3:
=-60000-6000/1.1-9000/1.1^2-14000/1.1^3+25500/1.1^3=-64252.4417731029

Year 4:
=-60000-6000/1.1-9000/1.1^2-14000/1.1^3-20000/1.1^4+19500/1.1^4=-83752.4759237757

Economic life is 1 year